OpenStudy (unklerhaukus):

f(x) =3sin(2x) + 2cos(2x)

OpenStudy (unklerhaukus):

\[=√(2^2+3^2)cos(2x- \phi); \phi = arctan(3/2)\]\[=√13cos(2x-\phi)\]

OpenStudy (unklerhaukus):

what is arctan(3/2)?

OpenStudy (anonymous):

\[0.983...\]OR \[56.31°\]

OpenStudy (unklerhaukus):

so the best answer is just \[=√13cos(2x−arctan(3/2))\]

OpenStudy (anonymous):

Is that \[=\sqrt{13\cos(2x−\arctan(3/2))}\]where the whole thing is rooted?

OpenStudy (unklerhaukus):

no, only the 13 is in the square root \[=\sqrt(13) cos(2x−arctan(3/2))\]

OpenStudy (anonymous):

Yes that is correct, but it is not neat, here is a more neat answer \[=√(13)\sin(2x+\arctan(2/3))\]

OpenStudy (unklerhaukus):

that is much better thankyou